A real estate investment has the following expected cash flows:

YEAR CASH FLOW
0 -$113,435.00
1 $12,051.00
2 $29,277.00
3 $50,339.00
4 $47,576.00


The investor wants a 6.00% return on this investment. What is the NPV of this opportunity?
Answer #1

DS fc =SUM(D3:07) 1 2 year Cash flows 3 0 $ 4 1 $ 2 $ 3 $ 4 $ pv@6% Present value (1,13,435.00) 1.0000 $ (1,13,435.00) 12,051

fax =SUM(D3:07) B pv@6% DS А 1 2 year 30 4 1 5 2 63 74 8 9 10 Cash flows -113435 12051 29277 50339 47576 =C3/1.06 =C4/1.06 =C